# DoRA: Weight-Decomposed Low-Rank Adaptation

### v4

@@ 2 Related Works
Parameter-Efficient Fine-Tuning (PEFT) methods are designed to reduce the high expense of fine-tuning large-scale models. They achieve this by training a relatively small subset of parameters, compared to the total number of parameters, for adapting to downstream tasks. Existing PEFT methods can be divided into three categories. The first category is referred to as Adapter-based methods, which involve introducing additional trainable modules into the original frozen backbone, such as (Houlsby et al., 2019; He et al., 2021; Karimi Mahabadi et al., 2021; mahabadi et al., 2021; Xia et al., 2024). For example, (Houlsby et al., 2019) proposes adding linear modules in sequence to the existing layer, whereas (He et al., 2021) advocates for integrating these modules in parallel with the original layer to enhance performance. The second category is Prompt-based methods. These methods add extra soft tokens (prompts) to the initial input and focus solely on fine-tuning these trainable vectors, as seen in works like (Lester et al., 2021; Razdaibiedina et al., 2023; Wang et al., 2023). However, these approaches typically face challenges due to their sensitivity to initialization, affecting their overall effectiveness. These first two categories, whether altering the model’s input or architecture, result in increased inference latency compared to the baseline model.
@@ 6.1 QDoRA: Enhancements to QLoRA
Figure 5: Accuracy comparison of LLaMA2-7B with QDoRA, QLoRA and FT on Orca-Math (Mitra et al., 2024).
@@
They conducted experiments on fine-tuning LLaMA2-7B using the Orca-Math(Mitra et al., 2024) dataset with QDoRA, QLoRA, and FT. The training set included 100k samples, with 500 reserved for evaluation using the exact match score as the metric. In addition to the fine-tuned models, they also reported results from zero-shot, few-shot, and FT with post-training quantization (PTQ), where the FT model is quantized to the BnB NF4 format after training. According to Figure 5, QDoRA not only significantly surpasses QLoRA by 0.19, but it also slightly outperforms FT by 0.05, while using considerably less memory. This indicates that QDoRA can effectively combines the parameter efficiency of QLoRA with the more granular optimization of full finetuning. These initial findings suggest that QDoRA holds considerable promise and could hugely benefit the opensoure community by substantially lowering the GPU memory requirements for fine-tuning large language models.
@@ 7 Conclusion
In this work, we first conduct a novel weight decomposition analysis to reveal the distinct learning patterns between LoRA and FT. Building on these insights, we introduce DoRA, a fine-tuning method that is compatible with LoRA and its variants and exhibits a closer resemblance to FT’s learning behavior. DoRA consistently outperforms LoRA across various fine-tuning tasks and model architectures. Specifically, DoRA improves upon LoRA in commonsense reasoning and visual instruction tuning tasks. Furthermore, DoRA also shows compatibility with VeRA on the Alpaca instruction tuning task. Moreover, DoRA can be considered as a costless alternative to LoRA, as its decomposed magnitude and direction components can be merged back into the pre-trained weight after the training, ensuring that there is no extra inference overhead. For future work, we wish to explore the generalizability of DoRA in domains beyond language and vision, particularly in the field of audio. Additionally, it would be intriguing to investigate the potential of DoRA in various other applications, such as Stable Diffusion text-to-image fine-tuning.
8 Acknowledgements

### v5

@@ 2 Related Works
Parameter-Efficient Fine-Tuning (PEFT) methods are designed to reduce the high expense of fine-tuning large-scale models. They achieve this by training a relatively small subset of parameters, compared to the total number of parameters, for adapting to downstream tasks. Existing PEFT methods can be divided into three categories. The first category is referred to as Adapter-based methods, which involve introducing additional trainable modules into the original frozen backbone, such as (Houlsby et al., 2019; He et al., 2021; Karimi Mahabadi et al., 2021; mahabadi et al., 2021). For example, (Houlsby et al., 2019) proposes adding linear modules in sequence to the existing layer, whereas (He et al., 2021) advocates for integrating these modules in parallel with the original layer to enhance performance. The second category is Prompt-based methods. These methods add extra soft tokens (prompts) to the initial input and focus solely on fine-tuning these trainable vectors, as seen in works like (Lester et al., 2021; Razdaibiedina et al., 2023; Wang et al., 2023). However, these approaches typically face challenges due to their sensitivity to initialization, affecting their overall effectiveness. These first two categories, whether altering the model’s input or architecture, result in increased inference latency compared to the baseline model.
@@ 6.1 QDoRA: Enhancements to QLoRA
Figure 5: Accuracy comparison of LLaMA2-7B/LLaMA3-8B with QDoRA, QLoRA and FT on Orca-Math (Mitra et al., 2024).
@@
They conducted experiments on fine-tuning LLaMA2-7B/LLaMA3-8B using the Orca-Math(Mitra et al., 2024) dataset with QDoRA, QLoRA, and FT. The training set included 100k samples, with 500 reserved for evaluation using the exact match score as the metric. In addition to the fine-tuned models, they also reported results from zero-shot, few-shot, and FT with post-training quantization (PTQ), where the FT model is quantized to the BnB NF4 format after training. According to Figure 5, QDoRA not only significantly surpasses QLoRA by 0.19/0.23 on LLaMA2-7B and LLaMA3-8B, but it also slightly outperforms FT on both models, while using considerably less memory. This indicates that QDoRA can effectively combines the parameter efficiency of QLoRA with the more granular optimization of full finetuning. These initial findings suggest that QDoRA holds considerable promise and could hugely benefit the opensoure community by substantially lowering the GPU memory requirements for fine-tuning large language models.
6.2 Text-to-Image Generation
Recently, as diffusion models have expanded in size, LoRA has become a popular method for efficiently fine-tuning large stable diffusion models. In this section, we aim to explore whether DoRA’s advantages over LoRA extend to the task of text-to-image generation. We follow the training pipeline of DreamBooth (Ruiz et al., 2023) for fine-tuning SDXL (Podell et al., 2023), utilizing the advanced training scripts developed by HuggingFace. The hyperparameter settings for LoRA and DoRA are kept the same, and we fine-tune the model using two challenging datasets: 3D icons and Lego sets. The sample seeds for generating the images are kept the same for LoRA and DoRA for fair comparison. The generated images are shown in Figure 8 and 9 in the appendix. The results indicate that DoRA achieves significantly better personalization than LoRA when using the same training settings, and more accurately reflects the training targets. For example, in Figure 8, the first sub-figure of DoRA’s output features a unique round square around the image, which is a feature common to all the training targets. In contrast, this feature is absent in all the LoRA outputs. A similar observation could be found with the Lego training targets, where only the DoRA outputs consistently incorporate the Lego logo in the generated images.
@@ 7 Conclusion
In this work, we first conduct a novel weight decomposition analysis to reveal the distinct learning patterns between LoRA and FT. Building on these insights, we introduce DoRA, a fine-tuning method that is compatible with LoRA and its variants and exhibits a closer resemblance to FT’s learning behavior. DoRA consistently outperforms LoRA across various fine-tuning tasks and model architectures. Specifically, DoRA improves upon LoRA in commonsense reasoning and visual instruction tuning tasks. Furthermore, DoRA also shows compatibility with VeRA on the Alpaca instruction tuning task. Moreover, DoRA can be considered as a costless alternative to LoRA, as its decomposed magnitude and direction components can be merged back into the pre-trained weight after the training, ensuring that there is no extra inference overhead. For future work, we wish to explore the generalizability of DoRA in domains beyond language and vision, particularly in the field of audio.
Acknowledgements
@@ A.7 Robustness of DoRA towards different rank
A.8 Text-to-Image Generation
Figures 8 and 9 show the images produced by SDXL fine-tuned with  DoRAand LoRA via DreamBooth (Ruiz et al., 2023) personalization techniques on two distinct training sets: 3D Icon [^1] https://huggingface.co/datasets/linoyts/3d_icon and Lego [^2] https://huggingface.co/datasets/merve/lego_sets_latest. The results reveal that DoRA can achieve considerably better personalization than LoRA with identical training configurations, more closely matching the training target.
Figure 8: Images generated with SDXL finetuned with LoRA and DoRA on the 3D Icon training sets.
Figure 9: Images generated with SDXL finetuned with LoRA and DoRA on the Lego training sets.